OpenStudy (anonymous):

Evaluate cot -220 degrees

OpenStudy (anonymous):

\[\cot(x) = \frac{ \cos(x) }{ \sin(x) }\] so can you find the values for cos(-220) and sin(-220) and find the exact value.

OpenStudy (anonymous):

how do you find those values?

OpenStudy (anonymous):

Do this on your calculator.\[\cos ^{-1}(-220) \] and \[\sin^{-1}(-220)\]

OpenStudy (anonymous):

i don't have a scientific calculator

OpenStudy (anonymous):

Use your regular one... don't really need one.... if not use the computer calculator.